what is the [H+] (and its error) of a solution of HCl that is prepared by using a 2mL volumetric pipet (tolerance +/0.01mL) of standardized HCl(0.02050+-0.00005 mol/L) and diluting to the mark with ddH2O in a 10mL volumetric flask (Tolerance +/0.02mL)
Answer #1

Solution of HCl :

Molarity : 2 mL*0.02050 M / 10 mL = 0.0041 M

[H+] = 0.0041 M

Uncertainty /error :

1 .    2mL volumetric pipet (tolerance +/0.01mL)   :

       % error = 0.01 mL /2 mL *100 = 0.5 %

2.   standardized HCl (0.02050+-0.00005 mol/L)

     % error = 0.00005 M / 0.02050 M *100 = 0.24 %

3.     10mL volumetric flask (Tolerance +/0.02mL) :

      % error = 0.02 mL/ 10 mL *100 = 0.2 %

Combined error = ( 0.52 + 0.242 + 0.22) = 0.6 %

thus, [H+] = 0.0041 M +- 0.6 %

or    [H+] = 0.0041 M +- 0.00003 M